# Li He1, Pei-Yang Lin1, Peng-Hu Wei5, Guo-Guang Zhao5, Yun-Zhe Liu†,1,4 1

### v1

@@ Abstract
Invasive brain-computer interfaces have garnered significant attention due to their high performance. The current intracranial stereoElectroEncephaloGraphy (sEEG) foundation models typically build univariate representations based on a single channel. Some of them further use Transformer to model the relationship among channels. However, due to the locality and specificity of brain computation, their performance on more difficult tasks, e.g., speech decoding, which demands intricate processing in specific brain regions, is yet to be fully investigated. We hypothesize that building multi-variate representations within certain brain regions can better capture the specific neural processing. To explore this hypothesis, we collect a well-annotated Chinese word-reading sEEG dataset, targeting language-related brain networks, over 12 subjects. Leveraging this benchmark dataset, we developed the Du-IN [^1] Du-IN refers to the phonetic transcription of ”{CJK}UTF8bsmi讀音” (i.e., pronunciation) in Chinese. model that can extract contextual embeddings from specific brain regions through discrete codebook-guided mask modeling. Our model achieves SOTA performance on the downstream 61-word classification task, surpassing all baseline models. Model comparison and ablation analysis reveal that our design choices, including (i) multi-variate representation by fusing channels in vSMC and STG regions and (ii) self-supervision by discrete codebook-guided mask modeling, significantly contribute to these performances. Collectively, our approach, inspired by neuroscience findings, capitalizing on multi-variate neural representation from specific brain regions, is suitable for invasive brain modeling. It marks a promising neuro-inspired AI approach in BCI.
@@ 1 Introduction
Brain signals refer to the biometric information collected from the brain. Their patterns provide valuable insights toward understanding the physiological functions of the brain and the mechanism of related diseases, leading to various applications, including speech decoding cho2023neural ; duan2023dewave ; moses2021neuroprosthesis , sleep cognition research liu2022bstt ; zheng2023universal , neurological disorders detection jiang2024large ; zhang2024brant , and so on. Due to the high signal-noise ratio, invasive recording methods (e.g., stereoElectroEncephaloGraphy (sEEG), ElectroCorticoGraphy (ECoG)) usually reveal these underlying mechanisms better than non-invasive recording methods. Compared with ECoG, sEEG imposes less trauma on patients and provides more stereotactic information from specific brain regions. Although some researchers moses2021neuroprosthesis ; metzger2023high have recently built high-performance speech decoders based on ECoG, there are few attempts made to build speech decoders based on sEEG.
Modeling intracranial neural signals, especially sEEG, has drawn much research attention, but several issues remain unresolved. Current research on modeling intracranial neural signals is predominantly divided into two lines according to the basic modeling unit (i.e., a single channel or a group of channels). Some studies wang2023brainbert ; zhang2024brant embed a single channel to build univariate representations, focusing primarily on downstream tasks that involve single-channel prediction, e.g., epilepsy detection. However, they haven’t validated their methods on more challenging tasks requiring integrating multiple channels for prediction, e.g., speech decoding. Other studies angrick2021real ; feng2023high fuse a group of channels to build multi-variate representations, mostly relying on fully supervised methods that heavily rely on labeled data. Nonetheless, labeling data at scale in medical experiments is often impractical or costly, emphasizing the need to maximize label efficiency. To overcome these limitations, self-supervised pre-training followed by fine-tuning can leverage abundant unlabeled data, enhancing performance on various downstream tasks.
@@
The primary challenge in modeling intracranial neural signals lies in extracting meaningful tokens, requiring careful consideration of two key factors. (1) Temporal scale. Since intracranial neural signals have high temporal resolution and signal-noise ratio, these tokens must capture rapid dynamic changes in brain activity. (2) Spatial scale. Since the functions of different brain regions are relatively specific, these tokens should correctly capture the information of each brain region for further integration. To better assess how well different models capture the intricate processing within each brain region, we can evaluate these methods on tasks mainly involving a few brain regions.
@@
Since speech mainly involves specific brain regions related to vocal production, as demonstrated in Section 2.1, we utilize speech decoding tasks to evaluate which model can effectively extract information from specific brain regions. Due to the lack of an open-source sEEG language dataset, we collected a well-annotated Chinese word-reading sEEG dataset (vocal production), including 12 subjects, which makes up for the problem of missing sEEG recordings in language tasks. Inspired by neuroscientific findings, we systematically demonstrate the locality and specificity of brain computation and propose the Du-IN model to solve the abovementioned issues. Compared to other existing methods for modeling brain signals, Du-IN achieves SOTA performance on the 61-word classification task, demonstrating the effectiveness of our model in extracting meaningful tokens that can capture both the rapid changes and the precise state of specific brain regions. It marks a promising neuro-inspired AI approach saxe2021if ; richards2019deep in BCI.
@@
    A novel framework for sEEG speech decoding – Du-IN, which learns multi-variate contextual embeddings through discrete codebook-guided mask modeling.
@@ 2.2 Language Decoding in BCI
The other key is well-designed models with good representations. Existing work for brain-to-language representations can be classified into two categories: self-supervision or alignment with representation models pre-trained on other modalities (e.g., text, audio). BrainBERT wang2023brainbert learns general embeddings through self-supervised mask modeling. DeWave duan2023dewave introduces discrete codex encoding and aligns neural representations with text embeddings from BART lewis2019bart , thus enhancing the extraction of semantic processing-related information from EEG recordings. Metzger1 et al. metzger2023high align neural representations with acoustic embeddings to improve the extraction of vocal production-related information from ECoG recordings.
@@ 2.3 Self-supervised Learning in BCI
In recent years, self-supervised pre-training has made significant progress in natural language processing devlin2018bert ; radford2018improving ; brown2020language and computer vision bao2021beit ; he2022masked ; chen2020generative . However, its potential in BCI is far from being explored. BrainBERT (for sEEG) wang2023brainbert builds univariate representations based on a single channel and utilizes mask modeling to learn general representations. Brant (for sEEG) zhang2024brant and LaBraM (for EEG) jiang2024large also build univariate representations but further consider the spatial correlation among channels. BENDR (for EEG) kostas2021bendr takes the other way – building multi-variate representations by fusing all channels, and it uses contrastive learning to learn contextual representations. MMM (for EEG) yi2024learning also builds multi-variate representations but further considers the difference among brain regions (i.e., splitting channels into different groups).
All existing pre-train methods for sEEG primarily construct univariate representations and some of them further employ Transformer models to capture relationships among channels. However, unlike EEG pre-training methods, their effectiveness over multivariate representations has not been experimentally proven. Besides, there is no standard channel configuration for sEEG recordings, unlike EEG recordings, which makes modeling spatial relationships in sEEG more challenging.
@@ 3 Method
The overall architecture of Du-IN is illustrated in Figure 2, where the raw sEEG signals are fused across channels to build multi-variate representations and further encoded for downstream tasks.
@@ 3.2 Model Architecture
Figure 2: The overall architecture of Du-IN Encoder, i.e., Neural Transformer. Du-IN Encoder is used as an encoder in all Du-IN models (i.e., Du-IN VQ-VAE, Du-IN MAE, Du-IN CLS), see Appendix C for more details. In Du-IN VQ-VAE, this encoder is called as "Neural Tokenizer". In Du-IN MAE & Du-IN CLS, this encoder is called as "Neural Encoder".
We introduce the neural Transformer, a general architecture for sEEG speech decoding tasks that can deal with any input sEEG signals with arbitrary time length, as shown in Figure 2. The key operation for archiving this is segmenting the sEEG signals into patches, inspired by patch embeddings in images dosovitskiy2020image . For each sample $\mathcal{X}$, we use a $W$-length window without overlap to segment it into patches, obtaining $\mathcal{X}=\{\bm{x}_{i}\in\mathbb{R}^{C\times W}|i=1,...,N\}$, where $N=\lfloor\frac{T}{W}\rfloor$ is the number of patches.
@@
As each sEEG patch has multiple channels, it is vital to fuse different channels to extract meaningful features before patch-wise interaction by self-attention. We employ a spatial encoder, which consists of a linear projection and several convolution blocks, to encode each sEEG patch into a patch embedding. The linear projection transforms the raw sEEG signals into the hidden neural space, and its weights are utilized for subsequent analysis. The convolution block is composed of a 1-D convolution layer and a batch normalization layer ioffe2015batch . We denote the output patch embeddings from the spatial encoder as
@@
Finally, the sequence of embeddings will be directly fed into the Transformer encoder vaswani2017attention to get the final encoded =$\mathcal{E}=\{\bm{e}_{i}\in\mathbb{R}^{d}|i=1,...,N\}$. To make the training of the Transformer more stable and efficient, we incorporate some modifications dehghani2023scaling . We add layer normalization to the queries and keys before the dot-product attention mechanism, which avoids over-large values in attention logits:
@@
3.3 Neural Tokenizer Training
Prior to pre-training DuIR through mask modeling, we need to tokenize the sEEG patches into discrete tokens. We introduce vector-quantized neural signal regression, which is trained by reconstructing the original sEEG signals, as shown in Figure 3. The key components are the neural tokenizer, which encodes the raw sEEG samples into embeddings, and the neural regressor, which reconstructs the original sEEG signals. The idea is basically inspired by VQ-VAE van2017neural , which encodes images into discrete latent embeddings.
Figure 3: Overview of neural tokenizer training and neural encoder pretraining. (a). We train the "Neural Tokenizer" to discretize sEEG signals into discrete neural tokens by reconstructing the original sEEG signals. (b). During the pre-training, part of sEEG patches are masked while the objective is to predict masked tokens from visible patches.
Neural Tokenizer.
We define a neural codebook $\mathcal{C}=\{\bm{c}_{i}|i=1,...,N_{codex}\}\in\mathbb{R}^{N_{codex}\times d_{codex}}$, where $N_{codex}$ is the number of the discrete neural embeddings and $d_{codex}$ is the dimension of each embedding. Given a sEEG sample $\mathcal{X}$, the neural tokenizer (i.e., the neural transformer illustrated in Figure 2) first encodes it to embeddings =$\mathcal{E}=\{\bm{e}_{i}\in\mathbb{R}^{d}|i=1,...,N\}$. After that, we utilize a linear projection $\bm{\mathrm{z}}_{c}$ get the mapped embeddings $\bm{\mathrm{z}}_{c}(\mathcal{E})=\{\bm{\mathrm{z}}_{c}(\bm{e}_{i})\in\mathbb{R}^{d_{codex}}|i=1,...,N\}$ in the codebook space. Then, the codebook looks up the nearest neighbor of each embedding $\bm{\mathrm{z}}_{c}(\bm{e}_{i})$ in the neural codebook $\mathcal{C}$. This procedure can be formulated as
@@ 3.3 Neural Tokenizer Training
where $\ell_{2}$ represents $\ell_{2}$ normalization and $\bm{\mathrm{z}}_{q}(\bm{e}_{i})$ is the quantized vector after the quantizer. This is equivalent to finding the closest neural embedding by cosine similarity and such $\ell_{2}$ normalization improves the codebook utilization peng2022beitv2 .
Neural Regressor.
The neural regressor consists of a Transformer decoder and a stack of transposed convolution layers. Given a sequence of the vector-quantized embeddings $\mathcal{Z}=\{\bm{z}_{i}|i=1,...,N\}$, the neural regressor convert these discrete embeddings back into raw sEEG signals $\tilde{\mathcal{X}}=\{\tilde{\bm{x}}_{i}|i=1,...,N\}$. The mean squared error (MSE) loss is utilized to guide the regression. The total loss for training the neural tokenizer (i.e., Du-IN VQ-VAE model) is defined as:
@@
where $\bm{\mathrm{sg}}$ represents the stop-gradient operation, which is an identity at the forward pass and has zero gradients. To stabilize the codebook update, we use the exponential moving average strategy van2017neural .
@@ 3.4 Pre-training Du-IN
$\mathcal{E}_{m}=\{\bm{e}^{m}_{i}|i=1,...,N\},\quad\bm{e}^{m}_{i}=\delta(i\in\mathcal{M})\odot\bm{e}_{[M]}+(1-\delta(i\in\mathcal{M})\odot\bm{e}^{p}_{i},$ (6)
where $\delta(\cdot)$ is the indicator function. After that, the masked embeddings $\mathcal{E}_{m}$ will be added by temporal embeddings, and then fed into the Transformer encoder. The output embeddings $\mathcal{E}$ will be used to predict the corresponding neural tokens through a linear classifier:

### v3

@@ Abstract
Invasive brain-computer interfaces with Electrocorticography (ECoG) have shown promise for high-performance speech decoding in medical applications, but less damaging methods like intracranial stereo-electroencephalography (sEEG) remain underexplored. With rapid advances in representation learning, leveraging abundant recordings to enhance speech decoding is increasingly attractive. However, popular methods often pre-train temporal models based on brain-level tokens, overlooking that brain activities in different regions are highly desynchronized during tasks. Alternatively, they pre-train spatial-temporal models based on channel-level tokens but fail to evaluate them on challenging tasks like speech decoding, which requires intricate processing in specific language-related areas. To address this issue, we collected a well-annotated Chinese word-reading sEEG dataset targeting language-related brain networks from 12 subjects. Using this benchmark, we developed the Du-IN [^1] Du-IN refers to the phonetic transcription of ”{CJK}UTF8bsmi讀音” (i.e., pronunciation) in Chinese. model, which extracts contextual embeddings based on region-level tokens through discrete codex-guided mask modeling. Our model achieves state-of-the-art performance on the 61-word classification task, surpassing all baselines. Model comparisons and ablation studies reveal that our design choices, including (i) temporal modeling based on region-level tokens by utilizing 1D depthwise convolution to fuse channels in the ventral sensorimotor cortex (vSMC) and superior temporal gyrus (STG) and (ii) self-supervision through discrete codex-guided mask modeling, significantly contribute to this performance. Overall, our approach – inspired by neuroscience findings and capitalizing on region-level representations from specific brain regions – is suitable for invasive brain modeling and represents a promising neuro-inspired AI approach in brain-computer interfaces. Code and dataset are available at https://github.com/liulab-repository/Du-IN.
@@ 1 Introduction
Brain signals refer to the biometric information collected from the brain. Their patterns provide valuable insights toward understanding the physiological functions of the brain and the mechanism of related diseases, leading to various applications, including speech decoding cho2023neural ; duan2023dewave ; moses2021neuroprosthesis , sleep cognition research liu2022bstt ; zheng2023universal , neurological disorders detection jiang2024large ; zhang2024brant , and so on. Due to the high signal-noise ratio, invasive recording methods (e.g., stereoElectroEncephaloGraphy (sEEG), ElectroCorticoGraphy (ECoG)) usually reveal these underlying mechanisms better than non-invasive recording methods. Many previous works jo2024eeg ; duan2023dewave have shown that decoding speech from EEG signals is difficult, and the performance is limited. Compared with ECoG, sEEG imposes less trauma on patients and provides more stereotactic information from specific brain regions. Although some studies moses2021neuroprosthesis ; metzger2023high have recently shown promise for building high-performance speech decoders based on ECoG, there are few attempts made to explore the potential of sEEG-based speech decoding.
Modeling intracranial neural signals, especially sEEG, has gained significant attention, but several issues remain unresolved. Current research on modeling neural signals is divided into two lines based on the basic modeling units (e.g., channel-level tokens or group-level tokens [^2] The term ”group-level” includes ”brain-level” and ”region-level,” and is distinct from ”channel-level.”). Some studies zhang2024brant ; jiang2024large utilize shared embedding blocks to embed single channels into channel-level tokens, neglecting the specificity of brain computation caucheteux2023evidence ; then they adopt spatial-temporal integration to model spatial relationships among them, attempting to regain the precise state of the brain. However, these methods mainly focus on channel-level classification tasks, e.g., seizure detection, yet fail to validate them on more challenging group-level classification tasks, e.g., speech decoding. Other studies duan2023dewave ; feng2023high fuse all channels (across the brain) to build brain-level tokens, overlooking the brain’s desynchronization nature buzsaki2006rhythms ; then they adopt temporal modeling to capture the rapid process of brain dynamics. Besides, labeling data at scale in medical experiments is often impractical or costly, emphasizing the need to maximize label efficiency. Hence, developing an efficient pre-training framework that draws on prior neuroscience findings is highly appealing, as it can make the most of abundant unlabeled data.
@@
The primary challenge in modeling intracranial neural signals lies in extracting meaningful tokens, requiring careful consideration of two key factors. (1) Temporal scale. Since intracranial neural signals have high temporal resolution and signal-noise ratio, these tokens must capture rapid dynamic changes in brain activity. (2) Spatial scale. Considering the brain’s desynchronization nature, these tokens should correctly capture the information of each brain region for further integration and, if needed, decouple different parts of brain dynamics within each brain region. To better assess how well different models capture the intricate processing within each brain region, we can evaluate these methods on tasks mainly involving a few brain regions.
@@
Since speech mainly involves specific brain regions related to vocal production, as demonstrated in Section 2.1, we utilize speech decoding tasks to evaluate which model can effectively extract information from specific brain regions. Since there are too few open source sEEG language datasets angrick2021real ; wang2023brainbert , we collected a well-annotated Chinese word-reading sEEG dataset (vocal production), including 12 subjects, which makes up for the problem of missing sEEG recordings in language tasks. Inspired by neuroscientific findings, we systematically demonstrate the locality and specificity of brain computation and propose the Du-IN model to solve the abovementioned issues. Compared to other existing methods for modeling brain signals, Du-IN achieves SOTA performance on the 61-word classification task, demonstrating the effectiveness of our model in extracting meaningful tokens that can capture both the rapid changes and the precise state of specific brain regions. It marks a promising neuro-inspired AI approach saxe2021if ; richards2019deep in BCI.
@@
    A novel framework for sEEG speech decoding – Du-IN, which learns region-level contextual embeddings through discrete codex-guided mask modeling.
@@ 2.2 Language Decoding in BCI
The other key is well-designed models with good representations. Existing work for brain-to-language representations can be classified into two categories: self-supervision or alignment with representation models pre-trained on other modalities (e.g., text, audio). BrainBERT wang2023brainbert learns general embeddings through self-supervised mask modeling. DeWave duan2023dewave introduces discrete codex encoding and aligns neural representations with text embeddings from BART lewis2019bart , thus enhancing the extraction of semantic processing-related information from EEG recordings. Metzger et al. metzger2023high align neural representations with acoustic embeddings to improve the extraction of vocal production-related information from ECoG recordings.
@@ 2.3 Self-supervised Learning in BCI
In recent years, self-supervised pre-training has made significant progress in natural language processing devlin2018bert ; radford2018improving ; brown2020language and computer vision bao2021beit ; he2022masked ; chen2020generative . However, its potential in BCI is far from being explored. BrainBERT (for sEEG) wang2023brainbert embeds single channels into channel-level tokens and utilizes mask modeling to learn general representations. Brant (for sEEG) zhang2024brant ; yuan2024brant , PopT (for sEEG) chau2024population and some works (for EEG) jiang2024large ; guetschel2024s further adopt spatial-temporal integration to model spatial relationships among them. Some works (for EEG) kostas2021bendr ; eldele2021time ; wu2024neuro ; foumani2024eeg2rep take the other way – fusing all channels (across the whole brain) to build brain-level tokens, and it uses self-supervised learning to learn contextual representations. Considering the difference among brain regions, MMM (for EEG) yi2024learning further splits channels into different groups to build region-level tokens.
All existing pre-training methods for sEEG primarily pre-train spatial-temporal models based on channel-level tokens yet only evaluate them on channel-level classification tasks, e.g., seizure detection. However, unlike EEG pre-training methods, their effectiveness over more challenging group-level classification tasks, e.g., speech decoding. Besides, there is no standard channel configuration for sEEG recordings, unlike EEG recordings, which makes modeling spatial relationships in sEEG more challenging.
@@ 3 Method
The overall architecture of Du-IN is illustrated in Figure 2, where the raw sEEG signals are fused across channels to build region-level tokens and further encoded for downstream tasks.
@@ 3.2 Model Architecture
Figure 2: The overall architecture of Du-IN Encoder. Du-IN Encoder is used as an encoder in all Du-IN models (i.e., Du-IN VQ-VAE, Du-IN MAE, Du-IN CLS (classification)), see Appendix C for more details.
We introduce the Du-IN Encoder, a general architecture for sEEG speech decoding tasks that can deal with any input sEEG signals with arbitrary time length, as shown in Figure 2. The key operation for archiving this is segmenting the sEEG signals into patches, inspired by patch embeddings in images dosovitskiy2020image . For each sample $\mathcal{X}$, we use a $W$-length window without overlap to segment it into patches, obtaining $\mathcal{X}=\{\bm{x}_{i}\in\mathbb{R}^{C\times W}|i=1,...,N\}$, where $N=\lfloor\frac{T}{W}\rfloor$ is the number of patches.
@@
As each sEEG patch has multiple channels, it is vital to fuse different channels to extract meaningful features before patch-wise interaction by self-attention. We employ a spatial encoder, which consists of a linear projection and several convolution blocks, to encode each sEEG patch into a patch embedding. The linear projection transforms the raw sEEG signals into the hidden neural space, and its weights are utilized for subsequent analysis. The convolution block is composed of a 1D depthwise convolution layer and a batch normalization layer ioffe2015batch . We denote the output patch embeddings from the spatial encoder as
@@
Finally, the sequence of embeddings will be directly fed into the Transformer encoder vaswani2017attention to get the final encoded =$\mathcal{E}=\{\bm{e}_{i}\in\mathbb{R}^{d}|i=1,...,N\}$. To make the training of the Transformer more stable and efficient, we incorporate some modifications dehghani2023scaling inspired by LaBraM jiang2024large . We add layer normalization to the queries and keys before the dot-product attention mechanism, which avoids over-large values in attention logits:
@@
3.3 Du-IN VQ-VAE Training
Prior to pre-training Du-IN through mask modeling, we need to tokenize the sEEG patches into discrete tokens. We introduce vector-quantized neural signal regression, which is trained by reconstructing the original sEEG signals, as shown in Figure 3. The key components are the Du-IN Encoder, which encodes the raw sEEG samples into embeddings, and the Du-IN Regressor, which reconstructs the original sEEG signals. The idea is basically inspired by VQ-VAE van2017neural , which encodes images into discrete latent embeddings.
Figure 3: Overview of Du-IN VQ-VAE training and Du-IN MAE training. (a). We train the Du-IN Encoder in the Du-IN VQ-VAE to discretize sEEG signals into discrete neural tokens by reconstructing the original sEEG signals. (b). During the training of Du-IN MAE, part of sEEG patches are masked while the objective is to predict masked tokens from visible patches.
Du-IN Encoder.
We define a neural codex $\mathcal{C}=\{\bm{c}_{j}|j=1,...,N_{codex}\}\in\mathbb{R}^{N_{codex}\times d_{codex}}$, where $N_{codex}$ is the number of the discrete neural embeddings and $d_{codex}$ is the dimension of each embedding. Given a sEEG sample $\mathcal{X}$, the Du-IN Encoder, illustrated in Figure 2, first encodes it to embeddings =$\mathcal{E}=\{\bm{e}_{i}\in\mathbb{R}^{d}|i=1,...,N\}$. After that, we utilize a linear projection $\bm{\mathrm{z}}_{c}$ to get the mapped embeddings $\bm{\mathrm{z}}_{c}(\mathcal{E})=\{\bm{\mathrm{z}}_{c}(\bm{e}_{i})\in\mathbb{R}^{d_{codex}}|i=1,...,N\}$ in the codex space. Then, the codex looks up the nearest neighbor of each embedding $\bm{\mathrm{z}}_{c}(\bm{e}_{i})$ in the neural codex $\mathcal{C}$. This procedure can be formulated as
@@ 3.3 Du-IN VQ-VAE Training
where $\ell_{2}$ represents $\ell_{2}$ normalization and $\bm{\mathrm{z}}_{q}(\bm{e}_{i})$ is the quantized vector after the quantizer. This is equivalent to finding the closest neural embedding by cosine similarity and such $\ell_{2}$ normalization improves the codex utilization peng2022beitv2 .
Du-IN Regressor.
The Du-IN Regressor consists of a Transformer decoder and a stack of transposed convolution layers. Given a sequence of the vector-quantized embeddings $\mathcal{Z}=\{\bm{z}_{i}|i=1,...,N\}$, the Du-IN Regressor convert these discrete embeddings back into raw sEEG signals $\tilde{\mathcal{X}}=\{\tilde{\bm{x}}_{i}|i=1,...,N\}$. The mean squared error (MSE) loss is utilized to guide the regression. The total loss for training the Du-IN VQ-VAE is defined as:
@@
where $\bm{\mathrm{sg}}$ represents the stop-gradient operation, which is an identity at the forward pass and has zero gradients. To stabilize the codex update, we use the exponential moving average strategy van2017neural .
@@ 3.4 Pre-training Du-IN
$\mathcal{E}_{m}=\{\bm{e}^{m}_{i}|i=1,...,N\},\quad\bm{e}^{m}_{i}=m_{i}\odot\bm{e}_{[M]}+(1-m_{i})\odot\bm{e}^{p}_{i},$ (6)
where $\delta(\cdot)$ is the indicator function and $m_{i}=\delta(i\in\mathcal{M})$. After that, the masked embeddings $\mathcal{E}_{m}$ will be added by temporal embeddings, and then fed into the Transformer encoder. The output embeddings $\mathcal{E}$ will be used to predict the indices of the corresponding codes from the codex in the Du-IN VQ-VAE through a linear classifier: